INFRARED RESONANCE ENHANCED PHOTODISSOCIATION SPECTROSCOPY OF METAL ION-BENZENE COMPLEXES

\underline{T.D. JAEGER} and M. A. DUNCAN, University of Georgia, Department of Chemistry, Athens, GA, 30602-2556; D. van HEIJNSBERGEN, G. von HELDEN, and G. MEIJER, F.O.M. Laboratory for Plasmaphysics, Rijnhuisen, Edinsonbaan 14, NL-3439 MN Nieuwegein, The Netherlands.

First row transition metal ion-benzene complexes of the form M^{+}(benzene)_{1}_{−}_{3} are produced by laser vaporization in a pulsed nozzle source, mass-selected in an ion-trap mass spectrometer and excited in the infrared with a free electron laser. Multiphoton dissociation occurs by the elimination of whole benzene molecules. The photofragment yield versus wavelength produces IR resonance-enhanced multiphoton photodissociation (IR-REMPD) spectra for these complexes. Vibrational bands observed in the 600-1800 cm^{−}^{1} region correspond to modes in the benzene molecules, but are shifted from those of free benzene due to the interaction with the metal ion.